OpenStudy (anonymous):

Can ANY1 HELP ME WITH MY MIDDLE SCHOOL HOME WORK PLEASE JUST 2 OR 1 QUESTIONS?????

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

question 1 @js777

OpenStudy (anonymous):

this one isnt upside down question 1 @js777

OpenStudy (texaschic101):

Its upside down..lol 1. 3/10 of 230 = 3/10 * 230 = 690/10 = 69 students voted for Luke 2. A = B*H B = 2 2/3 H = 1 5/6 turn them into improper fractions A = 8/3 * 11/6 A = 88/18 = 4 8/9 yds 3. 5 * 1/4 = 5/4 = 1 1/4 cups of butter

OpenStudy (anonymous):

can u help me wit more @texaschic101

OpenStudy (anonymous):

an TANK U @texaschic101

OpenStudy (anonymous):

can u help me wit more @texaschic101